OpenStudy (anonymous):

Perform the indicated operation and, if possible, simplify. Please show all of your work -4/9+3/10

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}=\frac{a d+b c}{b d} \]Use the formula above for adding two fractions together. You cross multiply and add the results to obtain the new numerator. Multiply the denominators together to get the new denominator. See how that works for you. The answer is\[\frac{-40+27}{90} \]

OpenStudy (anonymous):

thanx that help!